Jennifer's bakery sold half as many marble cakes as strawberry shortcakes. The price of a marble cakes is $6 and the price of a strawberry shortcake is $8.50. If the total amount of sales for these cakes was $391, what is the total number of each kind of cake sold? Please help?
Answer by stanbon(57397)

Jennifer's bakery sold half as many marble cakes as strawberry shortcakes. The price of a marble cakes is $6 and the price of a strawberry shortcake is $8.50. If the total amount of sales for these cakes was $391, what is the total number of each kind of cake sold?
---------------------
Equations:
Quantity: 2m = s
---
Value::: 6m + 8.5s = 391 dollars
----
Substitute for s and solve for "m":
6m + 8.5(2m) = 391
---
6m + 17m = 391
23m = 391
m = 17 (# of marble cakes)
---
solve for "s"
s = 2m
s = 34 (# of stawberry shortcake)
=====================================
Cheers,
Stan H.
=======